OpenStudy (anonymous):

How do you solve this equation? 24=13z-4z+6

OpenStudy (anonymous):

Combine like terms. 13z - 4z = 9z. Move 6 over 24 - 6 = 18 Divide 9 by both sides. 18/9 = 2 z = 2

OpenStudy (anonymous):

Okay.. Thank you!